OpenStudy (anonymous):

square root 27x^2y - square root 45x y^2 + square root 108x^2 y - square root 125x y ^2

OpenStudy (anonymous):

the options are 3x square root 15y-4y square root 5x x y square root 15x y square root 135x^2 y - square root 170 x y^2 9x square root 3y-8y square root 5x

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

haha im waiting..

OpenStudy (anonymous):

o.o

OpenStudy (anonymous):

?

OpenStudy (anonymous):

SORRY I HAD A CALL

OpenStudy (anonymous):

no worries haha

OpenStudy (anonymous):

soo.. help brittney time? :3

OpenStudy (anonymous):

CAN YOU PLEASE once again write the question clearly i am confused in square roots and powers

OpenStudy (anonymous):

\[\sqrt{27x ^{2}y}-\sqrt{45xy ^{2}}+\sqrt{108x ^{2}y}-\sqrt{125xy ^{2}}\] is this correct?

OpenStudy (anonymous):

\[\sqrt{27x ^{2}y} -\sqrt{45xy ^{2}} + \sqrt{108}x ^{2}y - \sqrt{125xy ^{2}}\]

OpenStudy (anonymous):

soo???

OpenStudy (anonymous):

\[x \sqrt{3*9y}+x \sqrt{3*36y}-y \sqrt{9*5x}-y \sqrt{25*5x}\] \[3x \sqrt{3y}+6x \sqrt{3y}-3y \sqrt{5x}-5y \sqrt{5x}\] \[9x \sqrt{3y}-8y \sqrt{5x}\]

OpenStudy (anonymous):

last option

OpenStudy (anonymous):

RIGHT

OpenStudy (anonymous):

Yes! thank you both very much!